OpenStudy (anonymous):

10sinθ^2+3sinθ-4=0 How many answers would i have? What are the solutions if 0360

OpenStudy (anonymous):

solving this gives two solutions and sin has already 2 so 4\[10a^2+3a-4=0\]

OpenStudy (anonymous):

Ok so my answers are +1/2 and -4/5

OpenStudy (anonymous):

that is my solution when i factor, so what are my solutions to find hte other 2?

OpenStudy (anonymous):

\[a=1/2,a=-4/5\] let \[\sin \theta=a\]

OpenStudy (anonymous):

Yes, so that is two. I am confused on how to find the others.

OpenStudy (anonymous):

\[\sin \theta=1/2\] \[\sin \theta =-4/5\] \[\theta=\sin ^{-1}*(1/2)+360k,k \in \mathbb{Z}\]

OpenStudy (anonymous):

and\[\theta=180-\sin ^{-1}(1/2)\]

OpenStudy (anonymous):

ah, and then 180- arcsin of -4/5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh wait, not 180, since it is not positive though

OpenStudy (anonymous):

arc sin of 4/5 is 53. So wouldnt you need the solution to be negative?

OpenStudy (anonymous):

\[180-\theta+360k\] k=1

OpenStudy (anonymous):

sorry, but i don't understand where the K is coming from

OpenStudy (anonymous):

its a general solution if you never did it then you can take it as positive

OpenStudy (anonymous):

So would you mind giving me the four solutions? I just want to know how to do this, but i may be understanding what you are saying wrong with the 180-theta+360k